OpenStudy (anonymous):

A cone has a diameter of 12 ft and a slant height of 20 ft. Explain whether tripling both dimensions would triple the surface area.

OpenStudy (welshfella):

multiplying a dimension by 3 will correspond to a 3^2 multiplication of the area.

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

hmmm how do you find the surface area of a cone anyway?

OpenStudy (anonymous):

S.A= pi(radius squared) + pi(radius x length)

OpenStudy (welshfella):

what is 3^2 ?

OpenStudy (anonymous):

9

OpenStudy (welshfella):

rightsurface area is multiplied by 9

OpenStudy (anonymous):

so no then?

OpenStudy (anonymous):

hmmm so... \(\textit{surface area of a cone}=\pi r{\color{brown}{ \sqrt{r^2+h^2}}}+\pi r^2\qquad \begin{cases} r=radius\\ {\color{brown}{ slant\ height}} \end{cases} \\ \quad \\ \textit{now multiplying both by 3, we get} \\ \quad \\ \pi (3)r{\color{brown}{ (3)\sqrt{r^2+h^2}}}+\pi (3)r^2\implies 9\pi r\sqrt{r^2+h^2}+3\pi r^2 \\ \quad \\ 3\left[ 3\pi r\sqrt{r^2+h^2}+\pi r^2 \right]\)

OpenStudy (anonymous):

a triple the surface area of 3SA would be \(\bf 3SA = 3\left[ \pi r\sqrt{r^2+h^2}+\pi r^2 \right]\ne 3\left[ 3\pi r\sqrt{r^2+h^2}+\pi r^2 \right]\)

OpenStudy (welshfella):

lets look at a simple example. This is applicable to all 3 dimensional shapes if we have a cube of side 2 cms S.A = 6*2^2 = 24 cm^2 if we triple the side lengths SA = 6* 6^2 = 216 the ratio of the area s is not 3:1 its 216 ; 24 = 9:1

OpenStudy (welshfella):

the ratio is the square of the ratio of side lengths

OpenStudy (welshfella):

the answer is NO